A piston cylinder device contains steam at 60 psia and 320F, and the temperature is increased to 400F through heat transfer. If the mass of the steam in the device is 15lbm and the weight of the piston is constant throughout the process. Calculate the work done during the process in Btu


1
Expert's answer
2021-05-10T02:42:40-0400

Steam in device = 60psia

Temperature = 320°

Increase in temperature = 400°

Mass of the steam = 15lbm


Solution:

The process of expansion is quasi-equilibrium

Thus,

"W_b = \int PdV = P_o \int dv = P_o (V_2-V_1)"

"W_b= mP_o(V_2 -V_1)"


According to the steam table,

V1 = 7.4863 and V2 = 8.3548


Substituting the value yields;

"W_b = (10) (60) (8.3548 - 7.4863)"

"( \dfrac1{ 5.404})= 96.4\ Btu"


"\therefore" The work done during the process is 96.4Btu.